OpenStudy (chels):

sin(2x)cos(5x) + cos(2x)sin(5x) use identities to simplify

myininaya (myininaya):

sin(7x)

OpenStudy (chels):

how did you get that?

myininaya (myininaya):

sin(2x+5x)

myininaya (myininaya):

remember sin(z+y)=sinzcosy+sinycosz where in this case z=2x and y=5x

OpenStudy (chels):

where did the cos go?

myininaya (myininaya):

its an identity

OpenStudy (chels):

this is so confusing.. :(

myininaya (myininaya):

do you know the identity sin(z+y)=sinzcosy+sinycosz

OpenStudy (chels):

yes

OpenStudy (chels):

but i don't understand how it's used in this question because sin and cos are being multiplied together.

myininaya (myininaya):

ok what if you let z=2x and y=5x what would you have?

OpenStudy (chels):

i understand how to use the identity, i'm just confused as to where it's being used

myininaya (myininaya):

ok we have [sin(2x)][cos(5x)]+[cos(2x)][sin(2x)]=[sin(z)][cos(y)]+[cos(y)][sin(z)]=sin(y+z) where y=2x and z=5x so this =sin(2x+5x)=sin(7x)

myininaya (myininaya):

oops i wrote the samething twice and that post

OpenStudy (chels):

ohhh my gosh. i get it now

OpenStudy (chels):

i'm so sorry. haha

myininaya (myininaya):

ok good i don't have to write it over correctly then?

OpenStudy (chels):

no, that is good 1 / 1-cos(x) + 1 / 1 + cos(x)

myininaya (myininaya):

i assume that (1-cosx) and (1+cosx) are in the denominator right?

OpenStudy (chels):

yes

myininaya (myininaya):

combine those fractions

OpenStudy (chels):

okay

myininaya (myininaya):

what do you get now?

myininaya (myininaya):

numerator=(1+cosx)+(1-cosx)=2 and denominator=1-(cosx)^2=(sinx)^2

myininaya (myininaya):

2/(sinx)^2=2(cscx)^2

OpenStudy (chels):

i got a denominator of 1-cos(x)1+cos(x)

myininaya (myininaya):

(1-cosx)(1+cosx) use foil here

OpenStudy (chels):

ohh okay

myininaya (myininaya):

or you could use the fact that (a-b)(a+b)=a^2-b^2

OpenStudy (chels):

we haven't learned that one, but i'll note it

myininaya (myininaya):

thats an algebra thing it should had been algebra you learned that not trig

OpenStudy (chels):

hmm, i don't remember it thankyou though.

myininaya (myininaya):

(x-3)(x+3)=x^2-9 and (x-5)(x+5)=x^2-25 and so on....

OpenStudy (chels):

\[\tan(\pi/5) + \tan(\pi/10) = numerator. 1 - \tan(\pi/5)\tan(\pi/10)\]

OpenStudy (chels):

sorry. let me re-write that

myininaya (myininaya):

what you got the one before this one?

OpenStudy (chels):

numerator = tan(pi/5) + tan(pi/10) denominator = 1 - tan(pi/5)tan(pi/10)

OpenStudy (chels):

yes, i understood it.

myininaya (myininaya):

Nice!

myininaya (myininaya):

Do you remember the idenity tan(x+y)=[tanx+tany]/[1-tanxtany]

OpenStudy (chels):

yes

myininaya (myininaya):

can use that here?

myininaya (myininaya):

x=pi/5 and y=pi/10

OpenStudy (chels):

3pi / 10?

OpenStudy (chels):

tan 3pi/10, i mean

myininaya (myininaya):

yes!!! :))

myininaya (myininaya):

ok if you want more help, i'm gonna be away from my computer so people will notice you you might want to start a new thread k? if you like me become a fan

OpenStudy (chels):

okay